### v2

# Reading Between the Lines: Towards Reliable Black-box LLM Fingerprinting via Zeroth-order Gradient Estimation

Abstract.
The substantial investment required to develop Large Language Models (LLMs) makes them valuable intellectual property, raising significant concerns about copyright protection. LLM fingerprinting has emerged as a key technique to address this, which aims to verify a model’s origin by extracting an intrinsic, unique signature (a “fingerprint”) and comparing it to that of a source model to identify illicit copies. However, existing black-box fingerprinting methods often fail to generate distinctive LLM fingerprints. This ineffectiveness arises because black-box methods typically rely on model outputs, which lose critical information about the model’s unique parameters due to the usage of non-linear functions. To address this, we first leverage Fisher Information Theory to formally demonstrate that the gradient of the model’s input is a more informative feature for fingerprinting than the output. Based on this insight, we propose ZeroPrint, a novel method that approximates these information-rich gradients in a black-box setting using zeroth-order estimation. ZeroPrint overcomes the challenge of applying this to discrete text by simulating input perturbations via semantic-preserving word substitutions. This operation allows ZeroPrint to estimate the model’s Jacobian matrix as a unique fingerprint. Experiments on the standard benchmark show ZeroPrint achieves a state-of-the-art effectiveness and robustness, significantly outperforming existing black-box methods.
Model Fingerprinting; AI Copyright Protection; Large Language Model; Trustworthy AI
journalyear: 2026conference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792196isbn: 979-8-4007-2307-0/2026/04ccs: Security and privacyccs: Computing methodologies Natural language processing
Resource Availability:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.18325480 and https://github.com/shaoshuo-ss/ZeroPrint.
1. Introduction
Large Language Models (LLMs) have become fundamental in revolutionizing the digital landscape (OpenAI, 2023; Guo et al., 2025), powering a vast array of web applications from intelligent search engines (Sancheti et al., 2024; Xiong et al., 2024) and conversational agents (Deng et al., 2024; Ning et al., 2025) to sophisticated content creation platforms (Comanici et al., 2025). However, the development of these state-of-the-art (SOTA) models demands substantial investment in computational resources, massive datasets, and extensive research expertise, rendering them highly valuable intellectual property for their developers (Wang and Kerschbaum, 2021; Zeng et al., 2024; Zhang et al., 2025a). Consequently, this value has given rise to significant security concerns regarding their copyright. Adversaries may engage in copyright infringement by using open-source models for commercial purposes without authorization (Shao et al., 2025a), or even resort to model stealing (Carlini et al., 2024; Yao et al., 2025), where proprietary models may be illicitly replicated and deployed. Protecting the copyright of these models is therefore not just a legal formality but a critical necessity to promote sustainable innovation and maintain a trustworthy AI ecosystem (Xu et al., 2025; Ren et al., 2024).
To address copyright concerns, LLM fingerprinting has emerged as a promising solution (Pasquini et al., 2025; Zeng et al., 2024; Zhang et al., 2025a; Gubri et al., 2024). This technique aims to determine whether a third-party suspicious model is derived from a copyrighted source model by extracting its intrinsic and unique features ($i.e.$, a fingerprint) in a non-invasive manner (McGovern et al., 2025; Zheng et al., 2022; Chen et al., 2022). Copyright attribution is then performed by measuring the similarity between the fingerprints of the two models. Compared to invasive LLM watermarking techniques (Li et al., 2023; Xu et al., 2024; Shao et al., 2025b), which proactively embed a verifiable signal into a model’s parameters or outputs before its release, fingerprinting offers significant advantages (Shao et al., 2025a). First, it does not require model fine-tuning, which means it has a smaller overhead and also does not degrade the model’s performance. Second, it can be applied to protect models that have already been deployed or released, a scenario where watermarking is not a viable option. Consequently, LLM fingerprinting has become a more practical and mainstream solution for copyright auditing.
Existing LLM fingerprinting techniques are primarily classified based on the level of access to the suspicious model, leading to two categories: white-box and black-box methods (Shao et al., 2025a). White-box fingerprinting assumes full access to a model’s internal components, including its architecture and parameters. This allows fingerprints to be extracted directly from static model weights (Zeng et al., 2024; Zheng et al., 2022), intermediate representations (Zhang et al., 2025a, 2024a), or gradients of weights (Wu et al., 2025). Conversely, black-box fingerprinting operates under the constraint that the model developer can only interact with the model through the API and observe the corresponding outputs. This paradigm is further divided into two strategies. The first, untargeted fingerprinting (Sun et al., 2025; Pasquini et al., 2025; Jin et al., 2024), involves submitting a selected set of queries to different models and comparing their outputs, where high similarity can indicate copyright infringement. The second strategy, targeted fingerprinting (Tsai et al., 2025; Gubri et al., 2024), generates a unique set of query-response pairs by optimizing them on a source model and then verifies if a suspicious model can consistently reproduce these specific outputs.
However, a significant performance gap persists between white-box and black-box fingerprinting methods. A recent study (Shao et al., 2025a) reveals that white-box techniques, especially static methods that extract features directly from model parameters, achieve near-perfect auditing performance. This effectiveness likely stems from the immense scale of LLM parameters, where the distinct training paths of independently developed models cause their static weights to become inherently unique identifiers. The success of static fingerprinting thus suggests that model parameters may be the most critical features for reliable auditing. A key challenge for black-box fingerprinting is therefore how to extract features that retain as much information about these parameters as possible. Existing black-box methods primarily rely on comparing model outputs. However, the extensive use of non-linear functions ($e.g.$, GELU and SiLU (Hendrycks and Gimpel, 2016)) throughout an LLM’s architecture progressively compresses parameter information, leading to significant information loss in the final output. This raises a critical research question for the field:
Is there another feature, accessible under black-box conditions, that can preserve more information about model parameters than output?
To address this question, we first ground our investigation in a formal analysis of feature information content based on Fisher Information Theory (Ly et al., 2017). Our theoretical analysis reveals a crucial insight: given the input $X$, the amount of information about the parameter $W$ that can be obtained by observing the gradient $D=\mathrm{d}Y/\mathrm{d}X$ is approximately always greater than that by observing the output $Y$. This finding suggests that the gradient is an inherently more distinctive feature for a reliable fingerprint. Based on this understanding, we propose ZeroPrint, a novel LLM fingerprinting method that captures this information-rich gradient in a black-box setting. The core challenge is that gradients are not directly accessible in the black-box auditing scenario. To overcome this, we adapt the principles of zeroth-order optimization (Chen et al., 2017), a technique designed to approximate gradients using only input-output queries. However, classical zeroth-order methods require applying small, continuous perturbations to inputs, which is a process not directly applicable to the discrete and symbolic nature of text. ZeroPrint tackles this gap by devising a method to simulate these perturbations in the text domain by generating semantically preserved variations of base queries through word substitution. By submitting these base and perturbed queries to a model, we can observe the corresponding changes in its output embeddings. These input-output difference pairs then allow us to estimate the local Jacobian matrix of the model using a regression-based approach, which serves as its unique fingerprint. By comparing the Jacobian fingerprint of a suspicious model with that of a source model, we can reliably determine its provenance.
Our contributions can be summarized as follows.
  • We provide a formal theoretical analysis grounded in Fisher Information Theory. Our analysis mathematically demonstrates that a model’s gradient with respect to its input contains more information about the model’s parameters than its output does, establishing gradients as a more robust feature for fingerprinting.
  • We introduce ZeroPrint, a novel black-box fingerprinting framework that estimates gradients as fingerprints without direct access to the model. ZeroPrint overcomes the challenge of operating in a discrete text domain by simulating input perturbations through semantic-preserving word substitutions, enabling the approximation of the Jacobian matrix as a distinctive fingerprint.
  • We conduct extensive experiments on LeaFBench, a representative benchmark for LLM copyright auditing. The results show that ZeroPrint consistently outperforms existing SOTA black-box fingerprinting methods across various metrics, demonstrating its superior effectiveness and reliability.
2. Background
2.1. Large Language Models
Large Language Models (LLMs) have become a cornerstone of the modern web, fundamentally reshaping how users interact with online services and information (Deng et al., 2024; Zhu et al., 2025; Luo et al., 2026). An LLM is a sophisticated computational system, typically defined by its vast parameter scale (Guo et al., 2025; Team, 2024c; Grattafiori et al., 2024). The generation of text by these models follows an auto-regressive paradigm. This means that the model produces text sequentially, one token at a time, where each new token is conditionally dependent on the sequence of previously generated tokens. The probability of generating a sequence of tokens $Y=(y_{1},y_{2},\dots,y_{T})$ is factorized into a product of conditional probabilities:
(1) $P(Y)=\Pi_{t=1}^{T}P(y_{t}|y_{1},y_{2},\dots,y_{t-1}).$
This underlying mechanism enables the diverse web applications, such as search engines (Sancheti et al., 2024), code generation (Nam et al., 2024; He et al., 2025), and web agents (Ning et al., 2025; Deng et al., 2024). LLMs represent one of the most significant advancements in computing and have become immensely valuable digital assets. Consequently, protecting the copyright of these models within the web environment has emerged as a critical challenge.
2.2. LLM Fingerprinting
As LLMs are integrated into a growing number of applications, protecting the intellectual property of these models has become a critical task (Xu et al., 2025; Li et al., 2025). LLM fingerprinting is a key technique that has emerged for this purpose, designed to verify a model’s provenance by analyzing its inherent characteristics (Shao et al., 2025a). In this process, a copyright auditor can extract a unique signature ($i.e.$, a fingerprint) from a suspicious model and compare it against the fingerprint of the source model to determine if it is an illicit copy or an unauthorized derivative. LLM fingerprinting provides a reliable mechanism for model owners to audit and protect their digital assets. Broadly, existing LLM fingerprinting can be categorized into two types: white-box and black-box LLM fingerprinting (Shao et al., 2025a).
White-box LLM Fingerprinting. White-box fingerprinting methods operate under the assumption that the auditor has full access to a model’s internal components, including its architecture and parameters. These techniques are categorized based on their feature source into three different methods: static, forward-pass, and backward-pass methods (Shao et al., 2025a). Static fingerprinting directly analyzes a model’s learned parameters as the primary identifier (Zeng et al., 2024; Zheng et al., 2022). Forward-pass methods leverage the dynamic intermediate representations generated as the model processes inputs (Zhang et al., 2025a, 2024a), while backward-pass techniques use the gradients produced during backpropagation as a unique signature (Wu et al., 2025). Although effective, the practical application of these methods is limited to open-source models where model parameters have already been accessible.
Black-box LLM Fingerprinting. In contrast, black-box fingerprinting methods are more broadly applicable, as they only require API-level access to query a model and observe its outputs. These approaches can be divided into two main categories: untargeted and targeted fingerprinting (Shao et al., 2025a).
  • Untargeted Fingerprinting: This is a straightforward and intuitive approach. The core idea is based on the observation that different LLMs often exhibit unique stylistic or behavioral “idiosyncrasies” in their generated outputs (Sun et al., 2025). Untargeted methods first select a set of queries, then submit them to different models and analyze the characteristics of the responses to generate a fingerprint (Pasquini et al., 2025; Gao et al., 2025; McGovern et al., 2025; Ren et al., 2025). A high degree of similarity between the fingerprints of a source and a suspicious model indicates that one is likely a derivative of the other.
  • Targeted Fingerprinting: This paradigm seeks to construct a specific set of query-response pairs that are unique to a source model and its derivatives (Gubri et al., 2024; Tsai et al., 2025; Jin et al., 2024). The key idea is that these special pairs act as a fingerprint, and only models from the source’s lineage will be able to consistently reproduce the predefined target responses. To perform an audit, the auditor submits these optimized queries to a suspicious model and checks if its outputs match the target responses. If a significant number of responses match, the model will be flagged as a likely derivative.
Despite the applicability of black-box methods, the empirical results from (Shao et al., 2025a) highlight a significant performance gap between the two paradigms. White-box methods demonstrate remarkable effectiveness and reliability due to their direct access to the model’s stable and unique parameter space. Conversely, black-box techniques are found to be less reliable, as they solely depend on the outputs that are inherently more stochastic and less stable. This discrepancy underscores that reliably inferring a model’s lineage with only black-box access remains challenging and highlights the significance of designing more effective solutions.
3. Motivation and Theoretical Analysis
The remarkable effectiveness of white-box fingerprinting, especially static methods that achieve near-perfect auditing performance by directly analyzing model parameters (Shao et al., 2025a), offers a crucial insight. This success suggests that model parameters may be the most reliable features for fingerprinting. The rationale is that the immense scale of LLM parameters causes the static weights of independently developed models to become inherently unique identifiers.
In the black-box scenario, therefore, the key challenge is to extract features that retain as much information about the model’s parameters as possible. In this section, we analyze two features observable under black-box conditions: the model output $Y$ and the gradient (or the differential) $D=\mathrm{d}Y/\mathrm{d}X$. Existing black-box methods primarily rely on the direct observation of model outputs. However, our theoretical analysis reveals that observing the gradient $D$ contains approximately more information about the parameter $W$ than the output $Y$. This finding indicates that $D$ is likely a more reliable feature for fingerprinting, forming the theoretical foundation for our ZeroPrint.
3.1. The Nonlinearity Bottleneck: A Motivation Example
To illustrate why observing the output $Y$ may yield less information about the model parameter $W$, let us consider a simple, fundamental unit commonly found in neural networks and LLMs:
(2) $Y=f(WX+K),$
where $X$ and $Y$ represent the input and output, respectively, while $W$ and $K$ are the weight and bias. The function $f$ is a non-linear activation function. An entire neural network, including LLM, is constructed by stacking a vast number of these non-linear units. In this example, the gradient of the output $Y$ with respect to the input $X$ can be expressed as:
(3) $D=\frac{\mathrm{d}Y}{\mathrm{d}X}=Wf^{\prime}(WX+K).$
By analyzing both $Y$ and $D$, a key difference emerges. If we observe the gradient $D$, for a fixed input $X$, the term $f^{\prime}(WX+K)$ becomes a constant value. Consequently, analyzing the characteristics of $D$ is effectively equivalent to analyzing the characteristics of the parameter $W$. In contrast, if we solely observe the output $Y$, the parameter $W$ must pass through the non-linear function $f$. Typically, irreversible non-linear functions ($e.g.$, GELU and SiLU (Hendrycks and Gimpel, 2016)) could compress the information, which can lead to a loss of information regarding $W$.
This observation leads us to a critical hypothesis: Is taking the gradient $D$ as a fingerprint a superior choice for preserving information about the parameter $W$ compared to taking the output $Y$? In the following section, we will analyze this hypothesis from the perspective of information theory.
3.2. The Theoretical Edge: Why Gradients Carry More Information
In this section, we provide a theoretical analysis of the hypothesis proposed in the preceding section. To formally quantify and compare the amount of information that an observable feature contains about a model parameter, we employ the concept of Fisher Information from information theory (Ly et al., 2017). Mathematically, Fisher Information quantifies how much an observable random variable $X$ carries about an unknown parameter $\theta$. A larger Fisher information value implies more information, which can in turn lead to a more precise estimate. We begin with the formal definition:
Definition 0 (Fisher Information).
Let $\{p_{\theta}(x)\}_{\theta\in\Theta}$ be a family of probability density functions parameterized by $\theta$. Assume $p_{\theta}(x)$ is differentiable with respect to $\theta$ and satisfies the regularity conditions that allow interchange of differentiation and integration. The Fisher information of a single observation $X$ about the parameter $\theta$ is defined as
(4) $\mathcal{I}_{X}(\theta):=\mathbb{E}_{X\sim p_{\theta}}\left[\left(\frac{\partial}{\partial\theta}\log p_{\theta}(X)\right)^{2}\right],$
where the expectation is taken with respect to $p_{\theta}(x)$.
Based on this definition, to analyze whether observation $D$ or observation $Y$ contains more information about the model parameter $W$, it is equivalent to comparing their respective Fisher Information with regard to the parameter $W$. Through this analysis, we arrive at the following theorem:
Theorem 2.
Let $Y=f(WX+K)$, where $f$ is neither a linear nor an affine function, $W$ and $K$ are constants with $W\neq 0$, and $D=\mathrm{d}Y/\mathrm{d}X=Wf^{\prime}(WX+K)$. Suppose that $X$ follows a zero-mean normal distribution, denoted as $X\sim\mathcal{N}(0,\sigma_{X}^{2})$, and $f^{\prime\prime}(K)\neq 0$. Under the first-order Taylor approximation, the Fisher information of $Y$ and $D$ with respect to $W$ satisfies:
(5) $\mathcal{I}_{D}(W)\geq(\frac{c_{1}^{2}}{2W^{2}c_{2}^{2}\sigma_{X}^{2}}+4)\cdot\mathcal{I}_{Y}(W),$
where $c_{1}=f^{\prime}(K),c_{2}=f^{\prime\prime}(K)$.
The proof is in Appendix A. Theorem 2 provides the mathematical proof for our hypothesis. The coefficient $c_{1}^{2}/(2W^{2}c_{2}^{2}\sigma_{X}^{2})+4>1$, leading to $\mathcal{I}_{D}(W)>\mathcal{I}_{Y}(W)$. The result formally establishes that the gradient $D$ contains approximately more information about the parameter $W$ than the output $Y$ does. This theoretical advantage confirms that the gradient is an inherently more distinctive and reliable feature for a robust model fingerprint.
4. Methodology
The theoretical analysis in the previous section provides strong evidence that gradients are more informative features for LLM fingerprinting than outputs. However, in a practical black-box scenario, we lack the direct access required for backpropagation to compute these gradients precisely. To bridge this gap, we introduce ZeroPrint, a novel black-box LLM fingerprinting method that leverages zeroth-order gradient estimation to approximate this information-rich feature. In the following sections, we will first define the threat model under which our method operates and then present a detailed introduction to the ZeroPrint framework.
4.1. Threat Model
In this paper, we consider a typical and strict threat model for black-box LLM fingerprinting, which involves two primary parties: the copyright auditor, who aims to protect their intellectual property, and a potential adversary, who seeks to infringe upon it.
Adversary’s Assumptions. The adversary’s primary motivation is to leverage a SOTA, copyrighted model to develop their own service with minimal investment. We assume the adversary obtains an illicit copy or an unauthorized derivative of a source model and deploys it for commercial use. The adversary possesses full control over this suspicious model and may perform modifications ($e.g.$, fine-tuning on a custom dataset or altering its behavior with different system prompts) to either adapt it for a specific task or attempt to obfuscate its origin.
Auditor’s Assumptions. The auditor operates under a set of constraints typical of a realistic black-box setting. We assume the auditor has the following capabilities:
  • The auditor has only black-box, API-level access to both the source and suspicious models. This access is limited to providing inputs and observing the corresponding textual outputs. Internal states and output logits are inaccessible in our threat model.
  • The auditor’s interactions with the source and suspicious LLMs are bound by a query limit to ensure practicality and mimic real-world API usage constraints. We assume the auditor can send a maximum of $q$ queries to a given LLM.
Figure 1. The pipeline of ZeroPrint. The process starts with creating a query set by generating semantically similar “Perturbed Queries” from a “Base Query” through word substitution. Both query types are fed into the LLM to get responses. All inputs and outputs are converted to embeddings to calculate input and output difference vectors. Finally, a linear regression model is fitted on these differences to estimate the Jacobian matrix, which serves as the model’s unique fingerprint.
4.2. Overview of ZeroPrint
Our ZeroPrint is structured as a three-step process, designed to effectively generate and compare model fingerprints in a black-box setting. The overall workflow is shown in Figure 1, as follows:
  • Step 1: Query Set Construction. The process begins with the construction of a query set, denoted by $\mathcal{Q}$. This set comprises $N$ input prompts $\{x_{1},x_{2},\dots,x_{N}\}$ that are designed to elicit distinctive responses from the models being audited.
  • Step 2: Fingerprint Generation. The query set $\mathcal{Q}$ is then fed into an LLM $\mathcal{M}$ to obtain a corresponding set of outputs, $Y$. Using these input-output pairs, we employ a zeroth-order gradient estimation technique to approximate the model’s Jacobian matrix for each input. The (aggregated) estimated Jacobian matrix, $J_{\mathcal{M}}$, serves as the unique fingerprint of the LLM. Given the source model $\mathcal{M}_{o}$ and the suspicious model $\mathcal{M}_{s}$, we extract their fingerprints $F_{o},F_{s}$ as follows:
    (6) $F=\text{Extract}(\mathcal{M}(\mathcal{Q}),\mathcal{Q})\triangleq J\approx\frac{\partial\mathcal{M}(\mathcal{Q})}{\partial\mathcal{Q}},$
    where $\mathcal{M}$ could be either $\mathcal{M}_{o}$ or $\mathcal{M}_{s}$.
  • Step 3: Fingerprint Comparison. Finally, the similarity between the fingerprint of the source model, $J_{o}$, and that of the suspicious model, $J_{s}$, is calculated to determine provenance. In our proposed ZeroPrint, we utilize the Pearson correlation coefficient, which can be interpreted as a form of normalized cosine similarity, as the primary metric. If the similarity score exceeds a predefined threshold $\tau$, the suspicious model is flagged as a derivative of the source model.
    (7) $\text{Sim}(J_{o},J_{s})=\text{Pearson}(J_{o},J_{s}).$
The core design of the ZeroPrint method hinges on two critical components: the construction of an effective query set and the specific algorithm used for fingerprint generation. A central challenge in this process is adapting gradient estimation for a black-box setting where inputs and outputs are discrete text.
In a conventional continuous domain, zeroth-order optimization estimates the gradient of a function $f:\mathbb{R}^{\zeta}\rightarrow\mathbb{R}^{\eta}$ by observing its output values. This is typically done by sampling a set of $m$ random direction vectors $\{u_{i}\}_{i=1}^{m}$ and querying the function at points perturbed along these directions, $f(x+\delta u_{i})$, where $\delta$ is a small smoothing parameter. The Jacobian matrix $J$ ($i.e.$, gradient) can then be approximated by solving a least-squares problem:
(8) $\min_{J}\sum_{i=1}^{m}\|(f(x+\delta u_{i})-f(x))-J(\delta u_{i})\|_{2}^{2}.$
Table 1. Overall performance comparison between ZeroPrint and baseline LLM fingerprinting methods.
White-box Black-box
Metric REEF LLMmap MET SEF TRAP ZeroPrint (Ours)
AUC $\uparrow$ 0.896($\pm$0.021) 0.632($\pm$0.009) 0.661($\pm$0.003) 0.581($\pm$0.008) 0.712($\pm$0.027) 0.720($\pm$0.013)
pAUC $\uparrow$ 0.832($\pm$0.019) 0.635($\pm$0.009) 0.658($\pm$0.003) 0.646($\pm$0.004) 0.625($\pm$0.006) 0.683($\pm$0.005)
TPR@1%FPR $\uparrow$ 0.634($\pm$0.021) 0.253($\pm$0.032) 0.329($\pm$0.004) 0.282($\pm$0.012) 0.227($\pm$0.020) 0.366($\pm$0.009)
MD $\uparrow$ 2.091($\pm$0.058) 0.521($\pm$0.021) 1.457($\pm$0.004) 0.172($\pm$0.014) 1.382($\pm$0.064) 1.457($\pm$0.015)
However, this approach is not directly applicable in our black-box scenario. The input and output spaces are discrete and symbolic ($i.e.$, text), not continuous vector spaces. It is not meaningful to add a small random vector to an input sentence. To overcome this, instead of adding continuous noise, ZeroPrint generates perturbed inputs by performing semantic-preserving word substitutions on the base queries. We elaborate on the details in the following subsections.
4.3. Query Set Construction
The process of constructing the query set $\mathcal{Q}$ is composed of two main substeps. First, we select $n$ base queries $\mathcal{Q}_{b}=\{x_{1},\dots,x_{n}\}$ designed for stability across a wide range of models. Second, for each base query $x_{i}$, we generate a corresponding set of $m$ perturbed queries $\mathcal{Q}_{p}^{(i)}=\{x^{\prime}_{i,1},\dots,x^{\prime}_{i,m}\}$ using a semantic-preserving word substitution technique, which is essential for the subsequent gradient estimation step. In total, we generate $n\times m+n=N$ samples as the Query Set $\mathcal{Q}$.
Base Query Selection. A key consideration in designing a generalizable fingerprinting method is the varying instruction-following capabilities across different LLM instances, especially between pre-trained base models and their fine-tuned derivatives. Base models, for instance, often struggle to produce coherent replies to complex queries, such as those used in QA tasks. To ensure our method elicits consistent and meaningful behavior from diverse LLMs, the selected queries could align with the fundamental objective of their pre-training, which is essentially a completion task. Therefore, we utilize a code completion task to form our base query set. This approach is broadly compatible with models regardless of their fine-tuning stage. Specifically, we extract code snippets from a standard code generation dataset ($e.g.$, HumanEval (Chen et al., 2021)) and prepend them with a simple instruction: “Complete the following code: [code snippet]”. From the dataset, we sample $n$ such code snippets to create our set of base queries.
Constructing Perturbed Queries via Word Substitution. With the set of base queries, the next step is to generate perturbed samples for each one, which will be used to estimate the Jacobian. For each base query, we generate $n_{p}$ perturbed counterparts using a word substitution method grounded in semantic similarity. The process leverages a pre-trained word embedding model ($e.g.$, GloVe (Pennington et al., 2014)) to find suitable replacements. For a given base query, we begin by randomly selecting $r$ replaceable words within the base query. For each of these $r$ words, we query the word embedding model to retrieve the top-$k$ most semantically similar words. A new, perturbed query is then formed by randomly replacing the original word with one of these $k$ candidates. This procedure is repeated independently to generate $m$ unique perturbed samples for each base query, resulting in a total of $n\times m$ perturbed queries. The example and the pseudocode can be found in Appendix B.
4.4. Fingerprint Generation via Zeroth-order Gradient Estimation
Once the query set is constructed, the next step is to generate the LLM’s fingerprint. This process involves three substeps: (1) querying the target LLM, (2) embedding the inputs and outputs into a continuous vector space, and (3) applying a regression method to estimate the Jacobian matrix, which serves as the final fingerprint.
Querying the Target LLM. The procedure begins by feeding the base queries $\{x_{i}\}_{i=1}^{n}$ and their corresponding perturbed counterparts $\{x^{\prime}_{i,j}\}_{j=1}^{m}$ to the target LLM, $\mathcal{M}$. To mitigate the inherent randomness of LLM outputs, each query is sent to the model $t$ times, yielding a set of textual responses.
Embedding Inputs and Outputs. We then employ a pre-trained sentence embedding model to transform all inputs and outputs into high-dimensional vectors. For each query, the final output embedding is computed by averaging the embeddings of the $t$ responses, resulting in a stable representation. Let $E(\cdot)$ be the sentence embedding function. The embedded vector for the base input $x_{i}$ is $e_{x_{i}}=E(x_{i})$, and its averaged output vector is $\bar{e}_{y_{i}}$. Similarly, the vectors for a perturbed query are $e_{x^{\prime}_{i,j}}=E(x^{\prime}_{i,j})$ and $\bar{e}_{y^{\prime}_{i,j}}$.
Estimate the Jacobian Matrix. With these vectors, we can formulate the Jacobian estimation as a linear regression problem. For each base query $x_{i}$, we compute the input and output difference vectors for all $m$ of its perturbations:
(9) Input difference: $\Delta x_{i,j}=e_{x^{\prime}_{i,j}}-e_{x_{i}},$
Output difference: $\Delta y_{i,j}=\bar{e}_{y^{\prime}_{i,j}}-\bar{e}_{y_{i}}.$
We seek to find a local Jacobian matrix $J_{i}$ that best models the linear relationship $\Delta y_{i,j}\approx J_{i}\Delta x_{i,j}$. Given the practical constraints on the number of queries allowed, the number of perturbation samples $m$ is often small relative to the dimensionality of the Jacobian matrix. This can lead to an underdetermined system, making an ordinary least squares estimate unstable. To address this, we use Ridge Regression, which introduces an L2 regularization term to produce a more robust estimate of the Jacobian:
(10) $J_{i}=\arg\min_{J}\sum_{j=1}^{m}\|\Delta y_{i,j}-J\Delta x_{i,j}\|_{2}^{2}+\alpha\|J\|_{F}^{2},$
where $\alpha$ is the regularization hyperparameter and $\|\cdot\|_{F}$ is the Frobenius norm. Finally, a Jacobian matrix $J_{i}$ is computed for each of the $n$ base queries. The overall fingerprint for the model $\mathcal{M}$, denoted as $J_{\mathcal{M}}$, is the element-wise average of these individual Jacobian matrices. This aggregation step creates a more comprehensive and stable fingerprint that captures the model’s characteristic behavior across a diverse set of inputs: $J_{\mathcal{M}}=\frac{1}{n}\sum_{i=1}^{n}J_{i}$.
With the final fingerprints generated for the source and suspicious models, the comparison is performed by calculating their similarity using Eq. (7) to determine copyright attribution.
5. Evaluation
5.1. Experimental Settings
Benchmark Models. The primary benchmark models in our experiments are from LeaFBench (Shao et al., 2025a). LeaFBench is built upon 7 mainstream pre-trained foundation models, including variants of Qwen (Team, 2024c), Llama (Grattafiori et al., 2024; Touvron et al., 2023), Mistral (Jiang et al., 2023), Gemma (Team, 2024a), and TinyLlama (Zhang et al., 2024b). These models were selected with parameter counts ranging from 1.1B to 14B. LeaFBench also implements various LLM post-development techniques, including fine-tuning, model merging, distillation, different system prompts, RAG, adversarial attacks, etc. In total, LeaFBench provides 149 different model instances and establishes a representative benchmark for evaluating fingerprint-based LLM copyright auditing. The details of LeaFBench can be found in Appendix B.
Baselines. For the baseline methods, we implement 3 untargeted black-box methods, LLMmap (Pasquini et al., 2025), Model Equality Testing (MET) (Gao et al., 2025), and Sentence Embedding Fingerprinting (SEF) (Shao et al., 2025a), and 1 targeted black-box method, TRAP (Gubri et al., 2024). We also include a white-box method, REEF (Zhang et al., 2025a), for reference. Note that targeted methods, including TRAP, are typically semi-black-box. They assume black-box access to the suspicious model, but white-box access to the source model, which is a slightly stronger assumption than untargeted methods. The maximum number of queries is set to $q=200$.
Settings of ZeroPrint. For default settings, we select $n=2$ base queries from HumanEval (Chen et al., 2021) and generate $m=4$ perturbed queries, with a total number of $10$ queries. We input each query to the LLM $t=20$ times, leading to $200$ queries in total. For the sentence embedding model, we choose all-mpnet-base-v2 (Reimers and Gurevych, 2019), which is a classic yet powerful model also used in LLMmap. More detailed settings are in Appendix B.
Metrics. In our experiments, following prior work (Shao et al., 2025a), we primarily leverage two metrics: (1) Area Under the ROC Curve (AUC) and (2) Partial AUC with FPR[$\in[0,0.05]$ (pAUC). The AUC provides a holistic measure of a method’s ability to distinguish between related model pairs (a source model and its derivative) and unrelated pairs across all possible classification thresholds $\tau$. In contrast, the pAUC metric is specifically designed to assess performance by focusing on a specific low false-positive-rate (FPR) region ($i.e.$, $[0,0.05]$) of the ROC curve. Higher AUC and pAUC scores indicate superior performance of a fingerprinting method.
Figure 2. The performance of ZeroPrint with different numbers of queries. We study three factors: the number of base queries $n$, the number of perturbed queries $m$, and the number of repeats per query $t$.
5.2. Main Results
Table 1 presents the overall performance comparison between ZeroPrint and various baseline methods. In addition to the primary metrics of AUC and pAUC, we also report the True Positive Rate at a 1% False Positive Rate (TPR@1%FPR) and the Mahalanobis Distance (MD). The TPR@1%FPR metric measures the true positive rate that a method can achieve while maintaining a low FPR of 1%. The MD metric measures the separability of the fingerprint distributions, reflecting the distance between positive pairs (a source model and its derivatives) and negative pairs (independent models). A larger MD indicates a stronger ability to distinguish between related and unrelated models. For both TPR@1%FPR and MD, higher values signify better performance.
The results demonstrate that ZeroPrint achieves SOTA or comparable performance across all evaluated metrics. It significantly outperforms existing untargeted methods. Notably, ZeroPrint also shows superior performance compared to the targeted, semi-black-box method TRAP, improving the AUC to 0.720 and the pAUC to 0.683. However, a performance gap remains when compared to white-box methods like REEF, which have direct access to model internals. It is worth noting that these results were achieved under a strict query limit of 200 queries per model. Subsequent experiments in our ablation study reveal that ZeroPrint’s performance can be further enhanced by increasing the number of queries allowed. This scalability is further detailed in Section 5.3.
5.3. Ablation Study
Effects of Different Numbers of Queries. The total number of queries is a critical factor influencing the performance of ZeroPrint. This number is determined by three key hyperparameters: the number of base queries ($n$), the number of perturbed queries generated for each base query ($m$), and the number of repeated generations for each query ($t$). To understand their impact, we conducted a series of experiments, varying one parameter at a time while keeping others fixed. The results are illustrated in Figure 2.
  • Effects of Different Numbers of Base Queries $n$. We first investigated the impact of $n$ by setting its value to 1, 2, 3, 4, and 5, with $n=2$ being the default in our main experiments. As shown in the first subplot of Figure 2, the performance of ZeroPrint remains relatively stable across different values of $n$. The AUC and pAUC fluctuate in a small interval. This observation suggests that the number of base queries may not have a significant impact on the effectiveness of ZeroPrint. A small number of base queries ($e.g.$, 2) is sufficient to generate a distinctive fingerprint.
  • Effects of Different Numbers of Perturbed Queries $m$. Next, we examined the influence of $m$ with values set to 1, 2, 4, 9, and 14. The default value used in our main experiments was $m=4$. The results, depicted in the second subplot of Figure 2, reveal that $m$ has a pronounced effect on the method’s performance. When $m$ is small ($e.g.$, 1 or 2), the estimated gradients are less stable, leading to inconsistent results ($e.g.$, a high AUC but a low pAUC). However, as m increases to 4 and beyond, both AUC and pAUC show a clear and significant upward trend. The performance peaks at $m=14$, achieving the highest AUC (0.778) and pAUC (0.706) in this set of experiments. This trend highlights the potential of ZeroPrint to achieve even greater effectiveness and robustness with an increased query budget, as more perturbed queries provide a more accurate gradient estimation.
  • Effects of Different Numbers of Repeats per Query $t$. Finally, we analyzed the effect of $t$, testing values of 1, 10, 20, 30, and 40, with $t=20$ as the default setting. This parameter is designed to mitigate the inherent randomness in LLM-generated outputs. The third subplot of Figure 2 shows that when $t$ is small, performance suffers due to this stochasticity, yielding an AUC of only 0.607 for $t=1$. As $t$ increases, the performance improves substantially, with the AUC reaching 0.720 at $t=20$. Beyond this point, the results begin to plateau and fluctuate slightly. This indicates that at $t=20$, the averaged output embedding for a given input becomes sufficiently stable. Further increases in $t$ offer diminishing returns.
Table 2. The performance of ZeroPrint using base queries from different datasets.
Dataset TruthfulQA SQuAD DROP Wikipedia HumanEval (Ours)
AUC $\uparrow$ 0.622 0.665 0.676 0.698 0.720
pAUC $\uparrow$ 0.668 0.666 0.651 0.647 0.683
Effects of Different Query Data. The choice of data for constructing base queries is another crucial element. We evaluated the performance of ZeroPrint using base queries derived from five different datasets: HumanEval (Chen et al., 2021) (our default), three question-answering (QA) datasets (TruthfulQA (Lin et al., 2022), SQuAD (Rajpurkar et al., 2016), and DROP (Dua et al., 2019)), and one text completion dataset (Wikipedia (Foundation, )). The results, presented in Table 2, show that the source of the query data significantly impacts fingerprinting effectiveness. When using queries from the QA datasets, the performance was consistently poor. Similarly, the Wikipedia dataset, despite being a completion task, also failed to achieve comparable results to HumanEval. This can be attributed to the nature of the tasks. Pure text completion is inherently more diverse and less structured than code completion. Thus, various derivative models tend to exhibit less consistent behavior on text-based prompts, making it harder to extract a stable and reliable fingerprint. In contrast, the structured nature of code completion elicits more uniform responses, leading to better performance.
Effects of Different Sentence Embedding Models. We investigated the influence of the sentence embedding model on ZeroPrint’s performance. We experimented with five different models of varying scales and embedding dimensions (details in Appendix B). The results, detailed in Table 3, reveal a clear trade-off related to the embedding dimension. When the dimension is too small, the model fails to adequately capture the semantic nuances of the sentences, leading to poor performance. Conversely, a larger embedding dimension increases the size of the Jacobian matrix to be estimated. Given the constraints on the number of queries, estimating this larger matrix becomes less accurate, which also degrades the overall fingerprinting effectiveness. This is supported by the similar performance of MPNet and EmbeddingGemma, which share the same embedding dimension and achieve comparable results.
Table 3. The performance of ZeroPrint using different sentence embedding models. “Dim” signifies the dimension of the generated sentence embeddings.
Dataset Model2Vec MiniLM MPNet (main) EmbeddingGemma Qwen3
Dim 256 384 768 768 1024
AUC $\uparrow$ 0.683 0.702 0.720 0.721 0.698
pAUC $\uparrow$ 0.659 0.682 0.683 0.682 0.659
5.4. Resistance to Potential Adaptive Attacks
Table 4. The fingerprint similarities between the original instruct LLM and its attacked version using ZeroPrint.
Attack Input Paraphrasing Output Perturbation
Setting 0.5B 1B 2B 0.05 0.10 0.15
Qwen2.5-7B 0.878 0.876 0.881 0.865 0.894 0.868
Llama3.1-8B 0.875 0.913 0.904 0.887 0.868 0.903
To assess the robustness of ZeroPrint, we evaluated its resilience against two potential adaptive attack strategies an adversary might employ to evade detection:
  • Input Paraphrasing Attack: This attack attempts to obfuscate queries by rewriting them using a separate, smaller LLM before submitting them to the target model. We simulated this using three paraphrasing models of varying sizes (0.5B, 1B, and 2B).
  • Output Perturbation Attack: This attack involves adding noise to the model’s output logits to disrupt the fingerprinting process. We tested this by introducing Gaussian noise with three different standard deviations (0.05, 0.10, and 0.15).
Detailed settings can be found in Appendix B. The experimental results, summarized in Table 4, demonstrate that ZeroPrint maintains high fingerprint similarity under both attack scenarios, confirming its strong resilience. The robustness against the input paraphrasing attack stems from our method’s reliance on the semantic difference between embeddings. While paraphrasing alters the surface form of the text, the core semantic relationship and vector difference between the base and perturbed queries remain relatively stable, allowing for a reliable estimation of the Jacobian matrix. For the output perturbation attack, ZeroPrint’s resilience can be attributed to two key factors. First, averaging the embeddings from multiple queries effectively filters out random, zero-mean noise. Second, since our fingerprint is derived from the estimated gradient, which is the relationship between input and output changes, it is inherently less sensitive to small, random perturbations in the absolute output values, thus preserving the fingerprint’s integrity.
Figure 3. The overhead of different fingerprinting methods.
5.5. Overhead Evaluation
To evaluate practical efficiency, we measured the total time for fingerprint extraction and verification on a subset of four LLMs. Experimental details can be found in Appendix B. From Figure 3, ZeroPrint demonstrates a competitive overhead of 424.9 seconds, which is significantly faster than most black-box baselines like MET, SEF, and TRAP. Its runtime is only greater than the white-box REEF and LLMmap, the latter of which uses a minimal set of only 8 queries. Given that the fingerprinting for a single model can be completed in under two minutes on average, ZeroPrint’s process completes in an acceptable timeframe. Overall, ZeroPrint is a computationally practical solution for real-world auditing.
6. Conclusion
In this paper, we addressed the critical challenge of protecting the intellectual property of large language models (LLMs) through fingerprinting, a technique that aims to identify a model’s origin by extracting a unique, intrinsic signature. We began by formally demonstrating, through Fisher Information Theory, that a model’s gradient contains significantly more information about its parameters than its output, establishing it as a superior feature for fingerprinting. Building on this theoretical foundation, we introduced ZeroPrint, a novel black-box fingerprinting method that successfully approximates these information-rich gradients. By simulating input perturbations through semantic-preserving word substitutions, ZeroPrint overcomes the inherent challenges of applying gradient estimation to the discrete domain of text. Our extensive experiments show that ZeroPrint not only achieves SOTA performance, outperforming existing black-box methods in both effectiveness and robustness, but also operates with a practical overhead. This work represents a significant step forward in developing reliable and secure mechanisms for LLM copyright auditing.
Acknowledgments
This research is supported in part by the “Pioneer” and “Leading Goose” R&D Program of Zhejiang (2024C01169), the Kunpeng-Ascend Science and Education Innovation Excellence/Incubation Center, and the National Natural Science Foundation of China under Grants (62441238, U2441240).
Appendix
Appendix A Proof of Theorem 2
Before proving Theorem 2, we first introduce and prove a related lemma about the Fisher information of a random variable following a normal distribution, as follows.
Lemma 0 (Fisher Information for a Normal Distribution).
Let $X$ be a random variable following a normal distribution with mean $\mu(\theta)$ and variance $\sigma^{2}(\theta)$, which are differentiable functions of a parameter $\theta$. The Fisher information of $X$ with respect to $\theta$ is:
(11) $\mathcal{I}_{X}(\theta)=\frac{1}{\sigma^{2}(\theta)}\left(\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}+\frac{1}{2\sigma^{4}(\theta)}\left(\frac{\partial\sigma^{2}(\theta)}{\partial\theta}\right)^{2}.$
Proof of Lemma 1.
The probability density function of the normal distribution is:
(12) $p(x;\theta)=\frac{1}{\sqrt{2\pi\sigma^{2}(\theta)}}\exp\left(-\frac{(x-\mu(\theta))^{2}}{2\sigma^{2}(\theta)}\right).$
The log-likelihood function is:
(13) $\log p(x;\theta)=-\frac{1}{2}\log(2\pi)-\frac{1}{2}\log(\sigma^{2}(\theta))-\frac{(x-\mu(\theta))^{2}}{2\sigma^{2}(\theta)}.$
The score function is the partial derivative of the log-likelihood with respect to $\theta$:
(14) $\frac{\partial}{\partial\theta}\log p(x;\theta)=\frac{(x-\mu(\theta))}{\sigma^{2}(\theta)}\frac{\partial\mu(\theta)}{\partial\theta}+\frac{(x-\mu(\theta))^{2}-\sigma^{2}(\theta)}{2\sigma^{4}(\theta)}\frac{\partial\sigma^{2}(\theta)}{\partial\theta}.$
The Fisher information is the expectation of the square of the score. Squaring the score function and taking the expectation, the cross-product term will have an expectation of 0. Thus, we only need to consider the expectation of the squared terms. Let $Z=X-\mu(\theta)\sim\mathcal{N}(0,\sigma^{2}(\theta))$, using the central moments of the normal distribution $\mathbb{E}[Z]=0$, $\mathbb{E}[(Z)^{2}]=\sigma^{2}(\theta)$, and $\mathbb{E}[(Z)^{4}]=3\sigma^{4}(\theta)$, we get:
(15) $\mathcal{I}_{X}(\theta)=\mathbb{E}\left[\left(\frac{(Z)}{\sigma^{2}(\theta)}\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}\right]+\mathbb{E}\left[\left(\frac{(Z)^{2}-\sigma^{2}(\theta)}{2\sigma^{4}(\theta)}\frac{\partial\sigma^{2}(\theta)}{\partial\theta}\right)^{2}\right].$
For the first term, we have:
(16) $\mathbb{E}\left[\left(\frac{(Z)}{\sigma^{2}(\theta)}\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}\right]$
$=$ $\frac{1}{\sigma^{4}(\theta)}\left(\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}\mathbb{E}[(Z)^{2}]=\frac{1}{\sigma^{2}(\theta)}\left(\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}.$
For the second term, we have:
(17) $\mathbb{E}\left[\left(\frac{(Z)^{2}-\sigma^{2}(\theta)}{2\sigma^{4}(\theta)}\frac{\partial\sigma^{2}(\theta)}{\partial\theta}\right)^{2}\right]$
$=$ $\frac{1}{4\sigma^{8}(\theta)}\left(\frac{\partial\sigma^{2}(\theta)}{\partial\theta}\right)^{2}\mathbb{E}[((Z)^{2}-\sigma^{2}(\theta))^{2}]$
$=$ $\frac{1}{4\sigma^{8}(\theta)}\left(\frac{\partial\sigma^{2}(\theta)}{\partial\theta}\right)^{2}(\mathbb{E}[(Z)^{4}]-2\sigma^{2}(\theta)\mathbb{E}[(Z)^{2}]+\sigma^{4}(\theta))$
$=$ $\frac{1}{4\sigma^{8}(\theta)}\left(\frac{\partial\sigma^{2}(\theta)}{\partial\theta}\right)^{2}(3\sigma^{4}(\theta)-2\sigma^{4}(\theta)+\sigma^{4}(\theta))$
$=$ $\frac{2\sigma^{4}(\theta)}{4\sigma^{8}(\theta)}\left(\frac{\partial\sigma^{2}(\theta)}{\partial\theta}\right)^{2}=\frac{1}{2\sigma^{4}(\theta)}\left(\frac{\partial\sigma^{2}(\theta)}{\partial\theta}\right)^{2}.$
Therefore, we get:
(18) $\mathcal{I}_{X}(\theta)=\frac{1}{\sigma^{2}(\theta)}\left(\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}+\frac{1}{2\sigma^{4}(\theta)}\left(\frac{\partial\sigma^{2}(\theta)}{\partial\theta}\right)^{2}.$
This completes the proof of Lemma 1. ∎
Given Lemma 1, we then provide the proof of Theorem 2.
Theorem 2.
Let $Y=f(WX+K)$, where $f$ is neither a linear nor an affine function, $W$ and $K$ are constants with $W\neq 0$, and $D=\mathrm{d}Y/\mathrm{d}X=Wf^{\prime}(WX+K)$. Suppose that $X$ follows a zero-mean normal distribution, denoted as $X\sim\mathcal{N}(0,\sigma_{X}^{2})$, and $f^{\prime\prime}(K)\neq 0$. Under the first-order Taylor approximation, the Fisher information of $Y$ and $D$ with respect to $W$ satisfies:
(19) $\mathcal{I}_{D}(W)\geq(\frac{c_{1}^{2}}{2W^{2}c_{2}^{2}\sigma_{X}^{2}}+4)\cdot\mathcal{I}_{Y}(W).$
where $c_{1}=f^{\prime}(K),c_{2}=f^{\prime\prime}(K)$.
Proof of Theorem 2.
We first analyze the Fisher information of $Y$ with respect to $W$. Let $Z=WX+K$. Since $X\sim\mathcal{N}(0,\sigma_{X}^{2})$, the random variable $Z$ is also normally distributed with mean $\mu_{Z}(W)=K$ and variance $\sigma_{Z}^{2}(W)=W^{2}\sigma_{X}^{2}$. We can now use Lemma 1 to calculate $I_{Z}(W)$ with the parameter $\theta=W$. The partial derivatives of the mean and variance of $Z$ with respect to $W$ are:
(20) $\frac{\partial\mu(W)}{\partial W}=0\quad\text{and}\quad\frac{\partial\sigma^{2}(W)}{\partial W}=2W\sigma_{X}^{2}.$
Substituting these into the formula from Lemma 1:
(21) $\mathcal{I}_{Z}(W)=\frac{1}{\sigma^{2}(W)}\left(\frac{\partial\mu(W)}{\partial W}\right)^{2}+\frac{1}{2\sigma^{4}(W)}\left(\frac{\partial\sigma^{2}(W)}{\partial W}\right)^{2}$
$=$ $\frac{1}{W^{2}\sigma_{X}^{2}}*0^{2}+\frac{1}{2(W^{2}\sigma_{X}^{2})^{2}}(2W\sigma_{X}^{2})^{2}=\frac{4W^{2}\sigma_{X}^{4}}{2W^{4}\sigma_{X}^{4}}=\frac{2}{W^{2}}.$
Since $Y=f(Z)$ and $f$ is independent of the parameter $W$, according to the Data Processing Inequality (DPI), the amount of information in data does not increase after any processing that is independent of the parameters. Therefore, we have:
(22) $\mathcal{I}_{Y}(W)\leq\mathcal{I}_{Z}(W)=\frac{2}{W^{2}}.$
Then, we analyze the Fisher information of $D=Wf^{\prime}(WX+K)$ with respect to $W$. Expand $f^{\prime}$ about $Z=K$:
(23) $f^{\prime}(Z)=c_{1}+c_{2}(Z-K)+R_{2}(Z),$
where $c_{1}:=f^{\prime}(K),c_{2}=f^{\prime\prime}(K)$, and $R_{2}(Z)$ is the Taylor remainder. Neglecting the remainder term $R_{2}$ (this is the first-order approximation), and using ZK=WXZ-K=WX, we obtain the approximate linear form as follows:
(24) $D\approx D_{\mathrm{approx}}:=c_{1}W+c_{2}W^{2}X.$
Under this approximation, $D_{\mathrm{approx}}$ follows a normal distribution $\mathcal{N}(\mu_{D}(W),\sigma^{2}_{D}(W))$ because $X$ also follows a normal distribution. From Eq. (24), we can calculate the mean and variance of $D_{\mathrm{approx}}$ as functions of $W$:
(25) $\mu_{D}(W)=\mathbb{E}[D_{\mathrm{approx}}]=Wc_{1},$
$\sigma^{2}_{D}(W)=\operatorname{Var}(D_{\mathrm{approx}})=W^{4}c_{2}^{2}\sigma_{X}^{2},$
since $\mathbb{E}[X]=0$ and $\operatorname{Var}(X)=\sigma_{X}^{2}$. Differentiate the mean and variance w.r.t. $W$:
(26) $\frac{\partial\mu_{D}(W)}{\partial W}$ $=c_{1},\ \ \frac{\partial\sigma^{2}_{D}(W)}{\partial W}=4W^{3}c_{2}^{2}\sigma_{X}^{2}.$
Invoking Lemma 1 with the parameter $\theta=W$, mean $\mu_{D}(W)$ and variance $\sigma^{2}_{D}(W)$, we obtain the Fisher information of the approximate normal model:
(27) $\mathcal{I}_{D}(W)$ $\approx\frac{1}{\sigma^{2}_{D}(W)}\Big(\frac{\partial\mu_{D}(W)}{\partial W}\Big)^{2}+\frac{1}{2\sigma^{4}_{D}(W)}\Big(\frac{\partial\sigma^{2}_{D}(W)}{\partial W}\Big)^{2}$
$=\frac{c_{1}^{2}}{W^{4}c_{2}^{2}\sigma_{X}^{2}}+\frac{1}{2\,W^{8}c_{2}^{4}\sigma_{X}^{4}}\,\Big(4W^{3}c_{2}^{2}\sigma_{X}^{2}\Big)^{2}$
$=\frac{c_{1}^{2}}{W^{4}c_{2}^{2}\sigma_{X}^{2}}+\frac{8}{W^{2}}=\frac{2}{W^{2}}(\frac{c_{1}^{2}}{2W^{2}c_{2}^{2}\sigma_{X}^{2}}+4).$
Therefore, we have the following conclusion: in this approximation,
(28) $\mathcal{I}_{D}(W)\geq(\frac{c_{1}^{2}}{2W^{2}c_{2}^{2}\sigma_{X}^{2}}+4)\cdot\mathcal{I}_{Y}(W).$
This completes the proof of Theorem 2. ∎
Base Query:
Complete the following code: def will_it_fly(q,w): ”’ Write a function that returns True if the object q will fly, and False otherwise. The object …
Perturbed Query:
1. Complete the following code : def will_it_fly ( q , w ) : “ ’ Write a function that returns Fact if the object q will flies , and False not . The object
2. Complete the following code : def will_it_fly ( q , w ) : “ ’ Book a function that returns Fact if the object q will fly , and False simply . The object
Base Query:
Complete the following code: def will_it_fly(q,w): ”’ Write a function that returns True if the object q will fly, and False otherwise. The object …
Perturbed Query:
1. Complete the following code : def will_it_fly ( q , w ) : “ ’ Write a function that returns Fact if the object q will flies , and False not . The object
2. Complete the following code : def will_it_fly ( q , w ) : “ ’ Book a function that returns Fact if the object q will fly , and False simply . The object
Example Queries in Query Set $\mathcal{Q}$
Figure 4. Example queries in the Query Set $\mathcal{Q}$.
Algorithm 1 ZeroPrint Query Set Construction
0:  $D_{source}$: Source dataset for base queries ($e.g.$, HumanEval)   $n$: Number of base queries to select   $m$: Number of perturbed queries to generate per base query   $r$: Number of words to replace in each perturbation   $k$: Number of top semantically similar candidates to consider   $E_{word}$: Pre-trained word embedding model ($e.g.$, GloVe)
0:  $Q$: The query set containing base and perturbed queries
1:  $Q_{b}\leftarrow\emptyset$ # Initialize set for base queries
2:  $Q\leftarrow\emptyset$ # Initialize the final query set
3:  for $i\leftarrow 1$ to $n$ do
4:   $x_{i}\leftarrow\text{SampleAndFormat}(D_{source})$ # e.g., “Complete the code: [snippet]”
5:   $Q_{b}\leftarrow Q_{b}\cup\{x_{i}\}$
6:  end for
7:  $Q\leftarrow Q\cup Q_{b}$ # Add all base queries to the final set
8:  for each base query $x_{i}\in Q_{b}$ do
9:   for $j\leftarrow 1$ to $m$ do
10:    $x^{\prime}_{i,j}\leftarrow x_{i}$ # Create a copy of the base query
11:    $W_{replaceable}\leftarrow\text{SelectRandomWords}(x_{i},r)$ # Select r words to replace
12:    for each word $w\in W_{replaceable}$ do
13:     $C_{candidates}\leftarrow E_{word}.\text{TopKSimilar}(w,k)$ # Find top-k similar words
14:     $w_{replacement}\leftarrow\text{RandomChoice}(C_{candidates})$
15:     $x^{\prime}_{i,j}\leftarrow\text{ReplaceWord}(x^{\prime}_{i,j},w,w_{replacement})$
16:    end for
17:    $Q\leftarrow Q\cup\{x^{\prime}_{i,j}\}$ # Add the perturbed query to the final set
18:   end for
19:  end for
20:  return $Q$
Algorithm 2 ZeroPrint Fingerprint Generation
0:  $\mathcal{M}$: The target black-box LLM   $Q_{b}=\{x_{i}\}_{i=1}^{n}$: Set of base queries   $Q_{p}=\{\{x^{\prime}_{i,j}\}_{j=1}^{m}\}_{i=1}^{n}$: Set of perturbed queries   $E$: Pre-trained sentence embedding model   $t$: Number of times to query the model for each input   $\alpha$: Regularization hyperparameter for Ridge Regression
0:  $J_{\mathcal{M}}$: The final fingerprint (aggregated Jacobian matrix)
1:  $J_{list}\leftarrow[]$ # Initialize list to store local Jacobians
2:  for each base query $x_{i}\in Q_{b}$ do
3:   $\Delta X_{i},\Delta Y_{i}\leftarrow[],[]$ # Initialize lists for input/output differences
4:   # Get stable embedding for the base query’s output
5:   $\{y_{i,k}\}_{k=1}^{t}\leftarrow\text{QueryModelMultipleTimes}(\mathcal{M},x_{i},t)$
6:   $\bar{e}_{y_{i}}\leftarrow\frac{1}{t}\sum_{k=1}^{t}E(y_{i,k})$
7:   $e_{x_{i}}\leftarrow E(x_{i})$
8:   for each perturbed query $x^{\prime}_{i,j}$ for $x_{i}$ do
9:    # Get stable embedding for the perturbed query’s output
10:    $\{y^{\prime}_{i,j,k}\}_{k=1}^{t}\leftarrow\text{QueryModelMultipleTimes}(\mathcal{M},x^{\prime}_{i,j},t)$
11:    $\bar{e}_{y^{\prime}_{i,j}}\leftarrow\frac{1}{t}\sum_{k=1}^{t}E(y^{\prime}_{i,j,k})$
12:    $e_{x^{\prime}_{i,j}}\leftarrow E(x^{\prime}_{i,j})$
13:    # Calculate input and output embedding differences
14:    $\Delta x_{i,j}\leftarrow e_{x^{\prime}_{i,j}}-e_{x_{i}}$
15:    $\Delta y_{i,j}\leftarrow\bar{e}_{y^{\prime}_{i,j}}-\bar{e}_{y_{i}}$
16:    Append $\Delta x_{i,j}$ to $\Delta X_{i}$
17:    Append $\Delta y_{i,j}$ to $\Delta Y_{i}$
18:   end for
19:   # Estimate the local Jacobian using Ridge Regression
[…]
21:   Append $J_{i}$ to $J_{list}$
22:  end for
23:  $J_{\mathcal{M}}\leftarrow\frac{1}{n}\sum_{i=1}^{n}J_{list}[i]$
24:  return $J_{\mathcal{M}}$
Appendix B Implementation Details
Introduction to LeaFBench. LeaFBench is a benchmark that facilitates a unified evaluation of LLM fingerprinting methods (Shao et al., 2025a). It includes 7 base model lineages, including Qwen2.5-7B, Qwen2.5-14B, Llama3.1-8B, Mistral-7B-v0.3, Gemma-2-2B, TinyLlama-1.1B-v1.0, and Llama-2-7B. LeaFBench also implements 13 different post-development techniques that may be adopted to adjust the pre-trained base models. These techniques can significantly affect the behavior of the LLMs. These techniques include: (1) pretraining, (2) instructional tuning, (3) fine-tuning, (4) parameter-efficient fine-tuning, (5) quantization, (6) model merging, (7) distillation, (8) general-purpose system prompts, (9) role-playing system prompts, (10) chain-of-thought prompts, (11) sampling strategies, (12) retrieval-augmented generation (RAG), (13) adversarial manipulation. LeaFBench includes a total number of 149 model instances.
Detailed Settings for ZeroPrint. To generate the base queries, we sample code snippets from HumanEval, take the first 20 words, and then prepend them with “Complete the following code:” to form a query for a completion task. When conducting word substitutions, we use GloVe (Pennington et al., 2014) as the word embedding model. For each base query, we randomly replace 3 words with one of the top-10 similar words. The example queries are shown in Figure 4 and the pseudocodes of query set construction and fingerprint generation in ZeroPrint are in Algorithm 1&2. The coefficient $\alpha$ in Ridge Regression is set to $0.001$. In our experiments, we utilize NVIDIA RTX 4090 GPUs with a total of 96GB VRAM.
Detailed Settings for the Baseline Fingerprinting Methods. In our experiments, we take 1 white-box method and 4 state-of-the-art black-box methods as our baseline. For fair comparison, following (Zhang et al., 2025a), we also set the query limit to $q=200$ while applying these methods. The detailed settings are as follows.
  • REEF (Zhang et al., 2025a): REEF is a forward-pass white-box fingerprinting method that relies on the intermediate representations as fingerprints. We strictly follow the original setting and sample 200 queries from TruthfulQA (Lin et al., 2022) as the input queries.
  • LLMmap (Pasquini et al., 2025): LLMmap is an untargeted black-box fingerprinting method. We adopt the 8 carefully-crafted prompts as the input queries and utilize the pre-trained open-set feature extraction model provided in its open-source repository.
  • MET (Gao et al., 2025): MET is also a untargeted black-box fingerprinting method that uses text completion prompts as input queries. The original MET uses 25 samples. Therefore, we run MET 8 times, leading to a total of 200 queries.
  • SEF (Shao et al., 2025a): SEF is an untargeted black-box fingerprinting method that straightforwardly uses the sentence embeddings of the outputs. Following original setting, we use Qwen3-Embedding-4B as the sentence embedding model and sample 200 queries from four different datasets, DROP (Dua et al., 2019), Ethics (Hendrycks et al., 2021), PubMedQA (Jin et al., 2019), and HumanEval (Chen et al., 2021).
  • TRAP (Gubri et al., 2024): TRAP is a targeted black-box fingerprinting method that leverages GCG (Zou et al., 2023) to generate targeted query-response pairs. For each query, we use GCG to optimize 100 iterations.
Detailed Settings of Ablation Studies. For the ablation study of different query data, except the HumanEval dataset (Chen et al., 2021), we select four other datasets as the source of the base queries, including the generative QA dataset TruthfulQA (Lin et al., 2022), the extractive QA dataset SQuAD (Rajpurkar et al., 2016), the complicated reasoning QA dataset DROP (Dua et al., 2019), and a text completion dataset Wikipedia (Foundation, ). For the Wikipedia dataset, similar to HumanEval, we sample a text snippet with 20 words and prepend a completion instruction ($i.e.$, “Continue the following text:”) to construct a completion prompt. For the ablation study of different sentence embedding models, we select the following 5 models: all-mpnet-base-v2 (MPNet), all-MiniLM-L6-V2 (MiniLM) (Reimers and Gurevych, 2019), potion-multilingual-128M (Model2Vec) (Tulkens and van Dongen, 2024), Qwen3-Embedding-0.6B (Qwen3) (Zhang et al., 2025b), and EmbeddingGemma-300m (EmbeddingGemma) (Schechter Vera et al., 2025). They are all classic or state-of-the-art sentence embedding models with different embedding dimensions.
Detailed Settings of Adaptive Attacks. For input paraphrasing attacks, we utilize three small LLMs: Qwen2-0.5B-Instruct (Team, 2024b), OLMo-2-0425-1B-Instruct (OLMo, 2024), and Gemma-2-2b-it (Team, 2024a), to rewrite the input queries. The experiments of adaptive attacks are conducted on Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct models.
Detailed Settings of the Overhead Evaluation. The overhead evaluation in Section 5.5 is conducted on four derivative models from the Qwen2.5-7B lineage. For fair comparison, we utilize 1 NVIDIA RTX 4090 48GB GPU to compute the runtime of all the evaluated LLM fingerprinting methods.
Table 5. The fingerprint similarities across different LLMs while auditing closed-source LLMs using ZeroPrint.
Type Deriative Independent Models
Models GPT-4o Gemini-2.5 Claude-4.5 DeepSeek-V3
GPT-4o 0.943 0.529 0.570 0.578
Appendix C Case Study of Closed-source Audit
In this section, we include a case study of auditing closed-source LLMs using ZeroPrint. Specifically, we utilize GPT-4o-mini (GPT-4o) as the source model and the same GPT-4o-mini API as the derivative, and take three other LLMs, Gemini-2.5-Flash-Lite (Gemini-2.5), Claude-Haiku-4.5 (Claude-4.5), and DeepSeek-V3, as independent models. The results in Table 5 show that the fingerprint similarity of fingerprints from the same GPT-4o-mini model is $0.943$, which is close to $1$, while other independent LLMs exhibit low similarity ($\approx 0.50$). These results demonstrate the feasibility and potential of our ZeroPrint to audit closed-source, black-box LLMs.
Appendix D Ethic Statement
The unauthorized usage of LLM has raised significant concerns about copyright infringements, while LLM fingerprinting offers a promising solution. In this paper, we propose a novel black-box fingerprinting method, ZeroPrint. Our method is designed for the purely defensive purpose of copyright auditing and does not introduce new threats or vulnerabilities. Furthermore, our work utilizes publicly available datasets, such as HumanEval, for query construction and does not infringe on the privacy of any individual. The research does not involve any human subjects. As such, this work does not raise any major ethical issues.